Troy Engines, Ltd., manufactures a variety of engines for use in heavy equipment. The company has always produced all of the necessary parts for its engines, including all of the carburetors. An outside supplier has offered to sell one type of carburetor to Troy Engines, Ltd., for a cost of $34 per unit. To evaluate this offer, Troy Engines, Ltd., has gathered the following information relating to its own cost of producing the carburetor internally: 19,000 Units Per Per Unit Year Direct materials Direct labor Variable manufacturing overhead Fixed manufacturing overhead, traceable Fixed manufacturing overhead, allocated Total cost s 16 s 304,000 10 190,000 38,000 9* 171,000 12 228,000 $ 49 ş 931,000 One-third supervisory salaries, two-thirds depreciation of special equipment (no resale value) Required: 1. Assuming the company has no alternative use for the facilities that are now being used to produce the carburetors, what would be the financial advantage (disadvantage) of buying 19,000 carburetors from the outside supplier? 2. Should the outside suppliers offer be accepted? 3. Suppose that if the carburetors were purchased, Troy Engines, Ltd., could use the freed capacity to launch a new product. The segment margin of the new product would be $190,000 per year. Given this new assumption, what would be the financial advantage (disadvantage) of buying 19,000 carburetors from the outside supplier? 4. Given the new assumption in requirement 3, should the outside suppliers offer be accepted? Complete this question by entering your answers in the tabs below. Required1 Required 2 Required 3Required 4 Assuming the company has no alternative use for the facilities that are now being used to produce the carburetors, what would be the financial advantage (disadvantage) of buying 19,000 carburetors from the outside supplier? Required 1 Required 2Required 1Required 2 Required 3 Required 4 Should the outside suppliers offer be accepted? Yes ONo K Required 1 Required 3 >Required 1Required 2 Required 3 Required 4 Suppose that if the carburetors were purchased, Troy Engines, Ltd., could use the freed capacity to launch a new product. The segment margin of the new product would be $190,000 per year. Given this new assumption, what would be the financial advantage (disadvantage) of buying 19,000 carburetors from the outside supplier? K Required 2 Required 4 >Required 1 Required 2 Required 3 Required 4 Given the new assumption in requirement 3, should the outside suppliers offer be accepted? Yes No Required 3 Required 4
Answer #1

1) Differential analysis

F

Make Buy
Direct material 304000
Direct labor 190000
variable manufacturing overhead 38000
Fixed manufacturing overhead (171000/3) 57000
Purchase cost (19000*34) 646000
Total relevant cost 589000 646000

Financial (disadvantage) = 589000-646000 = -57000

2) No

3) Differential analysis

Make Buy
Direct material 304000
Direct labor 190000
variable manufacturing overhead 38000
Fixed manufacturing overhead (171000/3) 57000
Opportunity Cost 190000
Purchase cost (19000*34) 646000
Total relevant cost 779000 646000

Financial advantage = 779000-646000 = $133000

4) Yes